OpenStudy (abbycross167):

Can someone please help me answer this math question? Please? I'll fan and give a medal!! I'm working on polynomial division Question - 30x + 6x --------- 6x

OpenStudy (gamer_girl2413):

i love division

OpenStudy (abbycross167):

Does that mean you know how to do this?

OpenStudy (gamer_girl2413):

I'll try what grade?

OpenStudy (abbycross167):

9th @gamer_girl2413

OpenStudy (photon336):

\[\frac{ 30x+6x }{ 6x }\] right?

OpenStudy (abbycross167):

yes sir/ma'am

OpenStudy (photon336):

so you notice both terms have an x in them right?

OpenStudy (photon336):

\[\frac{ 30x+6x = (36x) }{ (6x )}\]

OpenStudy (photon336):

now what can we do? \[\frac{ 36x }{ 6x } = ? \]

OpenStudy (abbycross167):

oh I was doing it wrong... I started by dividing 30 by 6

OpenStudy (abbycross167):

@Photon336 that would equal 6x

OpenStudy (photon336):